OpenStudy (anonymous):

triangle ABC has vertices A(–4, 4), B(6, 0), and C(–4, 0). Is triangle ABC a right triangle? @wolfe8

OpenStudy (gebooors):

Calculate lenghts of three sides, do they Realize Pyhtagoras equation?

OpenStudy (anonymous):

whats that?

OpenStudy (wolfe8):

An easier way is to draw it on a graph paper

OpenStudy (anonymous):

i don't have graphing paper, but would my answer be no?

OpenStudy (wolf1728):

I'll make the graphic

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

AC &BC make a right angle?

OpenStudy (wolf1728):

AC = 4 and BC = 10 If AC² + BC² = BA² then the triangle is a right triangle. 16 + 100 = BA² 116 = BA² BA = sqrt(116) = 10.7703296143 How do we know what the length of BA is anyway?

OpenStudy (anonymous):

so AC &BC do make a right angle, right?

OpenStudy (wolf1728):

I don't know (it sure looks it) How can we tell if angle C is a right angle?

OpenStudy (anonymous):

because the box in the graph you made, in fits it like a right angle

OpenStudy (wolf1728):

Yes, but mathematically, how can we prove it is a right angle? (Just looking like a right angle isn't enough).

OpenStudy (anonymous):

how do we prove it?

OpenStudy (whpalmer4):

We have two points at the same y value, no? If we can show that the third point connects to one of those points along a line perpendicular to the line connecting the first two points, we've demonstrated that there is a right angle.

OpenStudy (wolf1728):

How about calculating Angle B and angle A? If angle A + angle B = 90° then angle =90° arc tangent (angle B) = 4/10 = .4 angle B = 21.801° arc tangent (angle A) = 10/4 = 2.5 angle A = 68.199° angle A + anngle B = 90° Therefore angle C = 90°

OpenStudy (whpalmer4):

You can also show that angle C is a right angle directly. AC has undefined slope because the x-values are the same: the equation is x = constant. BC has slope 0 because the y-values are the same: the equation is y = constant. Such a pair of lines is perpendicular because they are parallel to the x and y axes, which are perpendicular.